Solve
s over 9 - 9 over s =0

Answer by Wolvstein(7)

s/9 - 9/s = 0
(Take 9/s to the other side)
s/9 = 9/s
s^2 = 81
s= 9
Check: 9/9 - 9/9 = 0
0 = 0